Displaying report 1-1 of 1.
Reports until 17:28, Thursday 05 December 2024
Controls General
Jeffrey Wack - posted 17:28, Thursday 05 December 2024 (12037)
Initial piezo lock of output filter cavity #2

[Jeff, Torrey]

We performed an initial lock of the second output filter cavity, and we eliminated the the 890Hz noise issue by putting the LNA10 Oscilloscope Preamplifier as an input buffer to the seeder DC modulation input.

Laser lock

The qualily of the lock with the laser is approximately as good as in B102. There is a 115kHz resonance with a high gain controller.

Piezo lock - no amplifier

We performed a lock with no amplifier by manually scanning the laser until we saw flashes, then activating the piezo loop. This loop was limited by a 2.6kHz resonance with a high gain controller

Piezo lock - with amplifier

The EVAL-ADHV4702-1 provides a gain of 21V/V, allowing us to scan on the piezo and lock. The 15V power supplies from Amazon are very bad and result in unacceptable 120Hz noise in the amplifier output. We switched to using a Newport 15V supply and this problem was eliminated. The lock with the amplifier is of a comparable quality to the laser lock, but it is limited by a 25kHz resonance. We will take data soon to better characterize the loop.

 

Displaying report 1-1 of 1.